# FocusChat: Text-guided Long Video Understanding via Spatiotemporal Information Filtering

### v1

@@ 4.4 Ablation Study
For few-shot ablation, each experiment in Tab. 5 involves pre-training and fine-tuning, with the same data used as in the previous zero-shot experiments. Since our model architecture is similar to Video-LLaMA, we use it as a reference for comparison. The “baseline” in Tab. 5 refers to replacing the trainable positional encoding in Video-LLaMA’s video Q-Former with our proposed absolute positional encoding VPE and adding an LN layer to each layer of the video Q-Former, as shown in Eq. 3. This improvement increased the model accuracy from 45.07 to 47.34. Rows from 2 to 6 in Tab. 5 show the impact of different alpha and beta parameter values in the STFM module on model performance. The optimal parameters were $\alpha$ = 1 and $\beta$ = 0, resulting in an accuracy of 48.93, an improvement of 1.59 over the baseline. This demonstrates that the effectiveness of the spatial-level feature filtering process aligning with prompt semantics. Rows of 7 and 8 in Tab. 5 introduce ablation of prompt-based temporal filtering (PBTF) module. With this addition, FocusChat’s accuracy increased to 49.4, validating the effectiveness of time-level feature filtering. Finally, to further confirm the effectiveness of VPE, we replaced the positional encoding in FocusChat with the original trainable positional encoding from Video-LLaMA. The accuracy dropped to 49.29, reaffirming the validity of the proposed positional encoding block.
@@ 5 Conclusion
In this paper, we introduce FocusChat, a text-guided model that employs spatiotemporal information filtering to realize the efficiency of visual information. As far as we know, we are the first to explore redundant spatiotemporal information filtering of visual features. The extracted visual tokens properly aligning with the user’s input elevate the model’s vision understanding capacity. Experiments confirm the effectiveness of FocusChat. In few-shot experiments, it obtained competitive performance on par with SOTA models, with only 10 percent of the pre-training data. It outperforms Video-LLaMA in zero-shot tasks with only 16 visual tokens used and a magnitude less data. Besides the research value, the proposed method will show its advantages in practical and constrained scenarios.

### v2

@@ 4.4 Ablation Study
For few-shot ablation, each experiment in Tab. 5 involves pre-training and fine-tuning, with the same data used as in the previous zero-shot experiments. Since our model architecture is similar to Video-LLaMA, we use it as a reference for comparison. The “baseline” in Tab. 5 refers to replacing the trainable positional encoding in Video-LLaMA’s video Q-Former with our proposed absolute positional encoding VPE and adding an LN layer to each layer of the video Q-Former, as shown in Eq. 4. This improvement increased the model accuracy from 45.07 to 47.34. Rows from 3 to 6 in Tab. 5 show the impact of different alpha and beta parameter values in the STFM module on model performance. The optimal parameters were $\alpha$ = 1 and $\beta$ = 0, resulting in an accuracy of 48.93, an improvement of 1.59 over the baseline. This demonstrates that the effectiveness of the spatial-level feature filtering process aligning with prompt semantics. Rows of 7 and 8 in Tab. 5 introduce ablation of prompt-based temporal filtering (PBTF) module. With this addition, FocusChat’s accuracy increased to 49.4, validating the effectiveness of time-level feature filtering. Finally, to further confirm the effectiveness of VPE, we replaced the positional encoding in FocusChat with the original trainable positional encoding from Video-LLaMA. The accuracy dropped to 49.29, reaffirming the validity of the proposed positional encoding block.
@@ 5 Conclusion
In this paper, we introduce FocusChat, a text-guided model that employs spatiotemporal information filtering to realize the efficiency of visual information. As far as we know, we are the first to explore redundant spatiotemporal information filtering of visual features by explicitly applying it to attention. The extracted visual tokens properly aligning with the user’s input elevate the model’s vision understanding capacity. Experiments confirm the effectiveness of FocusChat. In few-shot experiments, it obtained competitive performance on par with SOTA models, with only 10 percent of the pre-training data. It outperforms Video-LLaMA in zero-shot tasks with only 16 visual tokens used and a magnitude less data. Besides the research value, the proposed method will show its advantages in practical and constrained scenarios.